# Name Tagging Under Domain Shift via Metric Learning for Life Sciences

### v1

@@ 4.2 Main Results
Target Tasks
CHEMD BC5CDR DrugProt
Direct Transfer
46.18
78.55
68.17
Pse-Augment
47.14
76.43
68.20
Pse-Classifier
46.93
78.80
69.42
Ours
47.78 80.88 70.96
Table 4: F1(%) scores of our method compared to two alternative methods for using pseudo-labels.
@@ 4.3 Empirical Analysis
Pseudo Label Usage. Table 4 reports a comparison between our model and two alternative methods. Pse-Augment, where the detected pseudo-labels are marked and augmented with the target entities and a classifier trained to label unseen target entities. Pse-Classifier, where a classifier is trained to detect pseudo-labels and to filter them out, before being potentially mislabeled. This experiment aims to reveal the efficacy of the multi-similarity (MS) loss for discriminating between the source and the pseudo-labeled entities. The results indicate that our ED method that leverages MS loss is an effective way to use pseudo labels. For Pse-Augment, adding source domain entity labels to the target task leads to the negative transfer (NT) problem Zhang et al. (2020). For Pse-Classifier, the pipeline suffers from error propagation, where errors caused by the classifier can severely affect the performance of target entity predictions.
@@
How is the Representation Enhanced? To investigate the effect of our proposed methods, we project entity representations (i.e., averaged hidden states of entity tokens in input texts) into a two-dimensional space using t-SNE van der Maaten and Hinton (2008). We show the visualization of target task BC5CDR in Appendix F. For better comparison, we adopt the Davies-Bouldin index (DB) Davies and Bouldin (1979) as the criterion. The lower the DB is, the better the data points are clustered.
With our ED method, the model effectively learns to disperse the representations of chemical entities and pseudo labeled entities. Therefore, it becomes easier for our model to assign a negative label to a source domain entity by measuring similarities between its representations and the target domain entity representations. Furthermore, our EG method also plays a vital role in the formation of feature space of the target domain. In the projection, chemical entities are more clearly clustered and disease entities are dispersed from chemical entities, which makes chemical entities easier to be extracted. For a more precise comparison, Figure 3 reports the DB index. Our methods achieve lower DB compared to the baselines, which indicates that the improved representations of the target domain are learned with our methods.
@@ 5 Related Work
Entity extraction is a primary step in facilitating scientific discovery. Previous biomedical entity extraction methods can be categorized into several classes, including domain adaptive pretraining Labrak et al. (2023), boundary denoising diffusion model Shen et al. (2023a), question answering-based classification Arora and Park (2023), Cocke-Younger-Kasami (CYK) algorithm Corro (2023), in-context learning Chen et al. (2023), synthetic data Khandelwal et al. (2022); Chen et al. (2022); Hiebel et al. (2023), and prototype learning Cao et al. (2023).
Although there’s a shared corpus between the biomedical and chemical domains, entity extraction in the chemical domain remains underexplored. The chemical entity extraction task is usually viewed as an auxiliary task for biomedical named entity recognition Phan et al. (2021); Kocaman and Talby (2021); Luo et al. (2022); Lee et al. (2023). Similar to Nguyen et al. (2022), our paper differs from previous papers by viewing the chemical domain as an independent subject. Previous methods try to address this task by distant-supervision Wang et al. (2021b) or span-representation learning Nguyen et al. (2023). On the contrary, given the shared corpus between the biomedical and chemical domains, we leverage the large labeled data in the biomedical domain through transfer learning.
Transfer Learning for Name Tagging . Transfer learning is an effective method to address low-resource named entity recognition tasks Lee et al. (2018); Cao et al. (2018) and has shown its effectiveness in the biomedical domain Peng et al. (2019). Prior work has explored the role of continual pretraining on the target domain data Gururangan et al. (2020); Liu et al. (2021b). However, Mahapatra et al. (2022) argues that continual pretraining is inefficient regarding computational resources. In contrast to domain-adaptive pretraining, we aim to improve the representation of entities by projecting the source and target entities into separate regions of feature space. Inspired by the success of incorporating external knowledge for biomedical information extraction Zhang et al. (2021); Banerjee et al. (2021), we use biomedical events to augment the representations. To separate the potentially false positive examples in the target domain, we introduce pseudo-labels. Previous papers have adopted pseudo-labels in cross-lingual named entity recognition Zhou et al. (2023); Ma et al. (2023). However, they aim to align the entities in the source and target language instead of separating source entities from target entities. Compared to our method, Zhou et al. (2023) has a different contrastive objective, which aims to separate different entity types, rather than the entities from the two domains.
@@ 6 Conclusions
We proposed a named entity recognition task for transferring knowledge from the biomedical domain to the chemical domain. Our core idea is to train a shared feature space between the two domains to facilitate the knowledge transfer, and, at the same time, to project the source and target entities into separate regions of the feature space to reduce the false negative rate. We achieve this in a few steps. We begin by enriching the source feature space with information about events, then train a name tagging model to cluster similar entities into groups. We then use the trained model to label the entities that may belong to the source domain, and use these entities in a multi similarity loss function to achieve our goal. Our experiments across three source and three target datasets signify the effectiveness of our method.
@@ 7 Limitations
In this paper, we propose a framework that incorporates external knowledge during training. For instance, the compression function in §3.1 and templates in Section B.2 can be altered. Besides, such designs require prior knowledge of the source domain.
@@ Appendix A Details of ChatGPT evaluation on CHEMDNER
To evaluate ChatGPT’s few-shot ability of name tagging in the chemical domain, we elaborately select three name tagging examples in the CHEMDNER dataset that include all types of entities. An instructional description of the CHEMDNER task and examples constitute the prompt used to instruct the prediction. Configuration of prompts is shown in Table 5. The precision, recall, and macro-F1 scores on the CHEMDNER test set are $11.09\%$, $35.37\%$, and $16.88\%$ respectively. Laskar et al. (2023) reports the ChatGPT’s name tagging performance on general domain name tagging task-WNUT 17 dataset Derczynski et al. (2017) with precision: $18.03\%$, recall: $56.16\%$, and F1: $27.03\%$. This highlights ChatGPT’s weaker performance in solving chemical domain name tagging tasks.
Role
Prompt
System
You are an expert of chemical named entity recognition tasks
User
Description: In this task, you are given a small paragraph of a PubMed article, and your task is to identify all the named entities (particular chemical related entity) from the given input and also provide type of the each entity according to structure-associated chemical entity mention classes (ABBREVIATION, IDENTIFIER, FORMULA, SYSTEMATIC, MULTIPLE, TRIVIAL, FAMILY). Specifically, the paragraph are given with seperate tokens and you need to list all the chemical named entities in order and also tag their types. Generate the output in this format: entity1 <type_of_entity1>, entity2 <type_of_entity2>.
Examples:
Input: In situ C-C bond cleavage of vicinal diol following by the lactolisation resulted from separated treatment of Arjunolic acid ( 1 ) , 24-hydroxytormentic acid ( 2 ) and 3-O-$\beta$-D-glucopyranosylsitosterol ( 3 ) with sodium periodate and silica gel in dried THF according to the strategic position of hydroxyl functions in the molecule .
Output: C-C <FORMULA>, vicinal diol <FAMILY>, Arjunolic acid <TRIVIAL>, 24-hydroxytormentic acid <SYSTEMATIC>, 3-O-$\beta$-D-glucopyranosylsitosterol <SYSTEMATIC>, sodium periodate <SYSTEMATIC>, silica gel <TRIVIAL>, THF <ABBREVIATION>, hydroxyl <SYSTEMATIC>
Input: Structural studies using LC/MS/MS analysis and ( 1 ) H NMR spectroscopy showed the formation of a glycosidic bond between the primary hydroxyl group of RVX-208 and glucuronic acid .
Output: ( 1 ) H <FORMULA>, primary hydroxyl <FAMILY>, RVX-208 <IDENTIFIER>, glucuronic acid <TRIVIAL>
Input: The lystabactins are composed of serine ( Ser ) , asparagine ( Asn ) , two formylated/hydroxylated ornithines ( FOHOrn ) , dihydroxy benzoic acid ( Dhb ) , and a very unusual nonproteinogenic amino acid , 4,8-diamino-3-hydroxyoctanoic acid ( LySta ) .
Output: lystabactins <FAMILY>, serine <TRIVIAL>, Ser <FORMULA>, asparagine <TRIVIAL>, Asn <FORMULA>, formylated/hydroxylated ornithines <MULTIPLE>, FOHOrn <ABBREVIATION>, dihydroxy benzoic acid <SYSTEMATIC>, Dhb <ABBREVIATION>, 4,8-diamino-3-hydroxyoctanoic acid <SYSTEMATIC>, LySta <ABBREVIATION>
Please continue:
Input: %s
Output:
Table 5: Prompts of ChatGPT evaluation. %s is to be replaced by test context.
Appendix B Details of Event Embedding Construction.
B.1 Concatenation based Event Embedding
@@
B.2 Sentence-Encoder based Event Embedding
The prompt we use to instruct ChatGPT to generate explanatory templates for events is: give a one-sentence definition of biomedical event type XXX with arguments XXX, XXX…. For instance, we generate a template for the Phosphorylation event with give a one-sentence definition of biomedical event type Phosphorylation with arguments Trigger, Theme:Molecule, Cause:Molecule, Site:Simple chemical. The full list of the used templates is shown in Table 9, 10, 11, 12, and 13.
Appendix C Experiment Details
@@
Appendix D Full Experiment Results
More experiment results are reported in Table 8.
Target Tasks
CHEMD BC5CDR DrugProt
Direct Transfer
46.18
78.55
68.17
EG(MS)-Only
46.08
78.87
68.00
EG(concat)-Only
47.03
78.89
67.49
EG(sentEnc)-Only
46.91
79.76
67.93
Table 6: F1(%) scores on three target tasks. Performance of our EG method using vanilla MS loss without external knowledge is reported as EG(MS)-Only.
Appendix E Ablation Experiment Details
@@ E.1 Generalization Ability
Event Annotator
Target Tasks
CHEMD BC5CDR DrugProt
Gold-std
Concat
47.03
78.89
67.49
SentEnc
46.91 79.76
67.93
Auto-sys
Concat
46.06
79.49 68.31
SentEnc
46.11
79.39
68.04
Table 7: F1(%) scores of our proposed EG methods based on human/machine annotated events. We highlight better scores between Gold-std and Auto-sys annotators under each setting with underlines.
@@ Appendix G Scientific Artifacts
We list the license used in this paper: PathwayCuration (CC BY-SA 3.0), InfectiousDisease (CC BY-SA 3.0), CancerGenetics (CC BY-SA 3.0), CHEMDNER (CC BY 4.0), BC5CDR (CC BY 4.0), DrugProt (CC BY 4.0), Huggingface Transformers (Apache License 2.0), SapBERT (apache-2.0), S-PubMedBert-MS-MARCO-SCIFACT (apache-2.0), OpenAI (Terms of use666openai.com/policies/terms-of-use). We follow the intended use of all the mentioned existing artifacts in this paper.
Evaluate Datasets
CHEMDNER BC5CDR DrugProt
Metrics
Precision
Recall
F1
Precision
Recall
F1
Precision
Recall
F1
Few-shot (BERT)
42.73
51.69
46.77
72.44
85.86
78.51
63.80
67.42
65.49
@@
Direct Transfer
42.51
49.70
45.82
66.79
88.79
76.22
60.94
68.70
64.52
EG(concat)-Only
44.75
51.78
48.00
71.49
87.04
78.45
64.45
66.39
65.31
EG(sentEnc)-Only
42.80
50.43
46.25
70.91
86.97
78.09
64.36
67.81
66.03
ED-Only
46.97 52.30 49.48
74.71
83.96
79.06
67.05
67.13
67.08
EG(concat)+ED
43.44
51.54
47.11
72.34
86.09
78.61
66.59
66.13
66.32
EG(sentEnc)+ED
43.34
51.73
47.16
75.66
86.14
80.55
66.15
68.32
67.21
@@
Direct Transfer
41.57
48.30
44.65
74.92
82.51
78.53
61.96
67.81
64.75
EG(concat)-Only
45.26
50.65
47.77
72.28
87.41
79.12
63.76
65.25
64.43
EG(sentEnc)-Only
47.33
50.41
48.80
74.58
86.37
80.04
64.83
63.60
64.20
ED-Only
41.07
46.93
43.78
74.60
85.85
79.80
64.25
69.75
66.63
EG(concat)+ED
43.37
50.26
46.43
73.83
85.48
79.23
62.31
66.87
64.47
EG(sentEnc)+ED
41.50
52.19
46.18
74.70
85.95
79.93
65.06
69.11
67.02
@@
Direct Transfer
45.22
51.80
48.27
72.37
86.56
78.82
62.94
69.61
66.07
EG(concat)-Only
40.59
53.19
46.01
71.44
87.22
78.53
65.31
66.67
65.98
EG(sentEnc)-Only
41.54
52.36
46.33
72.68
86.09
78.82
66.99
65.71
66.30
ED-Only
47.06
50.58
48.75 75.08
86.96
80.57
66.26
73.52 69.68
EG(concat)+ED
45.16
51.72
48.16
73.37
85.06
78.76
65.60
67.67
66.64
EG(sentEnc)+ED
41.89
50.17
45.61
73.81
86.03
79.45
65.59
68.47
66.99
@@
Table 8: Full evaluation results. Pathway Curation, Infectious Diseases and Cancer Genetics are set to be the source domain respectively. Experiment settings are the same as main results reported in Table 3.
Type
Template
Conversion
@@
A specific trigger <Trigger> causes the transformation of a molecule <Theme> into another molecule <Product>.
Phosphorylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the addition of a phosphate group at a particular site <Site>, facilitated by another molecule <Cause>.
Dephosphorylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a phosphate group removed from a particular site <Site>, facilitated by another molecule <Cause>.
Acetylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of an acetyl group at a particular site <Site>, catalyzed by another molecule <Cause>.
Deacetylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has an acetyl group removed from a particular site <Site>, facilitated by another molecule <Cause>.
Ubiquitination
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the attachment of one or more ubiquitin molecules at a particular site, facilitated by another molecule <Cause>, often involving a simple chemical group <Site> as the site of attachment.
Deubiquitination
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has ubiquitin molecules removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
Hydroxylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of a hydroxyl group at a particular site, catalyzed by another molecule <Cause> involving a simple chemical group <Site> as the site of attachment.
Dehydroxylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a hydroxyl group removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
Methylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of a methyl group at a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of attachment.
Demethylation
@@
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a methyl group removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
Localization

### v2

@@ 4.2 Main Results
Target Tasks CHEMD BC5CDR DrugProt
Direct Transfer 46.18 78.55 68.17
Pse-Augment 47.13$\pm$0.8 76.43$\pm$0.2 68.20$\pm$0.8
Pse-Classifier 46.93$\pm$1.4 78.80$\pm$0.3 69.42$\pm$0.1
Ours 47.78$\pm$1.3 80.88$\pm$0.9 70.96$\pm$0.1
Table 4: F1(%) scores of our method compared to two alternative methods for using pseudo-labels. All the reported scores are averaged over 3 different random seeds. Additional experiment details are in Appendix E.3.
@@ 4.3 Empirical Analysis
Pseudo Label Usage.
Table 4 reports a comparison between our model and two alternative methods. Pse-Augment, where the detected pseudo-labels are marked and augmented with the target entities and a classifier trained to label unseen target entities. Pse-Classifier, where a classifier is trained to detect pseudo-labels and to filter them out, before being potentially mislabeled. This experiment aims to reveal the efficacy of the multi-similarity (MS) loss for discriminating between the source and the pseudo-labeled entities. The results indicate that our ED method that leverages MS loss is an effective way to use pseudo labels. For Pse-Augment, adding source domain entity labels to the target task leads to the negative transfer (NT) problem Zhang et al. (2020). For Pse-Classifier, the pipeline suffers from error propagation, where errors caused by the classifier can severely affect the performance of target entity predictions.
@@
How is the Representation Enhanced?
To investigate the effect of our proposed methods, we project entity representations (i.e., averaged hidden states of entity tokens in input texts) into a two-dimensional space using t-SNE van der Maaten and Hinton (2008) [^7] We show the visualization of target task BC5CDR in Appendix F.. For better comparison, we adopt the Davies-Bouldin index (DB) Davies and Bouldin (1979) as the criterion. The lower the DB, the better the clustering of the data points.
With our ED method, the model effectively learns to disperse the representations of chemical entities and pseudo-labeled entities. Therefore, it becomes easier for our model to assign a negative label to a source domain entity by measuring similarities between its representations and the target domain entity representations. Furthermore, our EG method also plays a vital role in the formation of feature space of the target domain. The projection results in clearer clustering of chemical entities, while disease entities are dispersed from them, making it easier to extract chemical entities. For a more precise comparison, Figure 3 reports the DB index. Our methods achieve lower DB compared to the baselines, which indicates that the improved representations of the target domain are learned with our methods.
Method Precision Recall F1
Target Only 41.32$\pm$1.9 45.52$\pm$2.0 43.31$\pm$1.9
Direct Transfer 42.16$\pm$1.3 48.51$\pm$1.4 45.08$\pm$0.6
EG(sentEnc) 45.49$\pm$1.0 49.03$\pm$1.9 47.16$\pm$0.8
ED 45.35$\pm$3.2 47.88$\pm$3.5 46.58$\pm$3.3
EG(sentEnc)+ED 44.94$\pm$0.4 49.37$\pm$2.0 47.03$\pm$0.8
Table 5: Averaged F1 scores (%) over 3 different random seeds for CHEMDNER trained on full BERT model.
Role of Adapters
To clarify that the use of adapters does not interfere with the conclusion of our proposed methods, we additionally finetune the full BERT model on the CHEMDNER dataset, and the results are reported in Table 5. The performance of the full model is similar to the performance of adapters or even slightly worse than our adapter models (47.03 vs 48.26). Our methods remain effective when fine-tuned with the full model, demonstrating that the results in the paper are reliable and sufficient.
Domain #Train/Valid/Test #Train/Valid/Test (All)
Science 28 / 66 / 543 200 / 450/ 543
AI 13 / 46 / 431 100 / 350 / 431
Table 6: Overview of CrossNER dataset.
Target Only Direct Transfer EG(sentEnc) ED
0.59 16.74 19.12 17.18
Table 7: Averaged F1 scores (%) over three different random seeds for the CrossNER dataset, transferring from the Science domain to the AI domain.
Compatibility Across Other Domain Pairs
In the above experiments, we focus on transfer learning between the biomedical domain and the chemical domain. To show the generalization ability of our proposed framework on other domain pairs, we conduct the experiments on two additional domains based on CrossNER Liu et al. (2021b), transferring from the Science domain to the AI domain. These two domains are highly related and share similar named entities. We downsample the train and validation data to roughly 10% of the full dataset for the target domain. Detailed statistics are shown in Table 6. The F1 scores averaged over three runs are reported in Table 7. It shows that our methods have strong generalization ability.
@@ 5 Related Work
Entity extraction is a primary step in facilitating scientific discovery Wang et al. (2021a). Previous biomedical entity extraction methods can be categorized into several classes, including domain adaptive pretraining Labrak et al. (2023), boundary denoising diffusion model Shen et al. (2023a), question answering-based classification Arora and Park (2023), Cocke-Younger-Kasami (CYK) algorithm Corro (2023), in-context learning Chen et al. (2023), synthetic data Khandelwal et al. (2022); Chen et al. (2022); Hiebel et al. (2023), and prototype learning Cao et al. (2023).
Although there’s a shared corpus between the biomedical and chemical domains, entity extraction in the chemical domain remains underexplored. The chemical entity extraction task is usually viewed as an auxiliary task for biomedical named entity recognition Phan et al. (2021); Kocaman and Talby (2021); Luo et al. (2022); Lee et al. (2023). Similar to Nguyen et al. (2022) and Wang et al. (2024), our paper differs from previous papers by viewing the chemical domain as an independent subject. Previous methods try to address this task by distant-supervision Wang et al. (2021c) or span-representation learning Nguyen et al. (2023). On the contrary, given the shared corpus between the biomedical and chemical domains, we leverage the large labeled data in the biomedical domain through transfer learning.
Transfer Learning for Named Entity Recognition.
Transfer learning is an effective method to address low-resource named entity recognition tasks Lee et al. (2018); Cao et al. (2018) and has shown its effectiveness in the biomedical domain Peng et al. (2019). Prior work has explored the role of continual pretraining on the target domain data Gururangan et al. (2020); Liu et al. (2021b). However, Mahapatra et al. (2022) argues that continual pretraining is inefficient regarding computational resources. In contrast to domain-adaptive pretraining, we aim to improve the representation of entities by projecting the source and target entities into separate regions of feature space. Inspired by the success of incorporating external knowledge for biomedical information extraction Zhang et al. (2021); Banerjee et al. (2021), we use biomedical events to augment the representations. To separate the potentially false positive examples in the target domain, we introduce pseudo-labels. Previous papers have adopted pseudo-labels in cross-lingual named entity recognition Zhou et al. (2023); Ma et al. (2023). However, they aim to align the entities in the source and target language instead of separating source entities from target entities. Compared to our method, Zhou et al. (2023) has a different contrastive objective, which aims to separate different entity types, rather than the entities from the two domains.
@@ 6 Conclusions
We proposed a named entity recognition task for transferring knowledge from the biomedical domain to the chemical domain. Our core idea is to train a shared feature space between the two domains to facilitate the knowledge transfer, and, at the same time, to project the source and target entities into separate regions of the feature space to reduce the false negative rate. We achieve this in a few steps. We begin by enriching the source feature space with information about events, then train a named entity recognition model to cluster similar entities into groups. We then use the trained model to label the entities that may belong to the source domain, and use these entities in a multi-similarity loss function to achieve our goal. Our experiments across three sources and three target datasets signify the effectiveness of our method.
@@ 7 Limitations
In this paper, we propose a framework that incorporates external knowledge during training. For instance, the compression function in §3.1 and templates in Section C.2 can be altered. Besides, such designs require prior knowledge of the source domain.
@@
Acknowledgements
This work is supported by U.S. DARPA ITM FA8650-23-C-7316, by the Molecule Maker Lab Institute: an AI research institute program supported by NSF under award No. 2019897, by DOE Center for Advanced Bioenergy and Bioproducts Innovation U.S. Department of Energy, Office of Science, Office of Biological and Environmental Research under Award Number DESC0018420, by U.S. the AI Research Institutes program by National Science Foundation and the Institute of Education Sciences, Department of Education through Award No. 2229873 - AI Institute for Transforming Education for Children with Speech and Language Processing Challenges, and by AI Agriculture: the Agriculture and Food Research Initiative (AFRI) grant no. 2020-67021- 32799/project accession no.1024178 from the USDA National Institute of Food and Agriculture. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied of, the National Science Foundation, the U.S. Department of Energy, and the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix A Details of ChatGPT evaluation on CHEMDNER
Figure 4: Components of concatenation based event embeddings. Arguments of events, along with event type, are encoded by an off-the-shelf model and concatenated afterwards. For nested events as arguments, we fill in compressed event embeddings recursively.
To evaluate ChatGPT’s few-shot ability of named entity recognition in the chemical domain, we elaborately select three named entity recognition examples in the CHEMDNER dataset that include all types of entities. An instructional description of the CHEMDNER task and examples constitute the prompt used to instruct the prediction. Configuration of prompts is shown in Table 8. The precision, recall, and macro-F1 scores on the CHEMDNER test set are $11.09\%$, $35.37\%$, and $16.88\%$ respectively. Laskar et al. (2023) reports the ChatGPT’s named entity recognition performance on general domain named entity recognition task-WNUT 17 dataset Derczynski et al. (2017) with precision: $18.03\%$, recall: $56.16\%$, and F1: $27.03\%$. This highlights ChatGPT’s weaker performance in solving chemical domain named entity recognition tasks.
Appendix B Notation Table
We present definitions for all notations we used in Table 9.
Notation Definition
$\mathbf{X}$ Set of input documents
$\mathbf{Y}$ Set of named entity tags
$n_{e}$ Total number of entities
$x_{i}$ The $i$-th entity
$y_{i}$ Label of the $i$-th entity
$f$ Classifier model
$\mathbb{S}$ Source domain
$\mathbb{T}$ Target domain
$S_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$ Relative similarity
$E(\mathord{\color[rgb]{0.67,0.67,0.67}\bullet})$ Set of embedding vectors
$\kappa_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$ Auxiliary similarity
$\mathcal{P}_{i}$ Set of selected positive entities
$\mathcal{N}_{i}$ Set of selected negative entities
$\mathcal{Q}$ Set of entities with pseudo labels
$I^{+}$, $J^{+}$ Auxiliary term for positive pairs
$I^{-}$, $J^{-}$ Auxiliary term for negative pairs
$w^{+}$ Soft weight of positive pairs
$w^{-}$ Soft weight of negative pairs
$\hat{w}_{ij}^{+}$ Adjusted soft weight of positive pairs
$\hat{w}_{ij}^{-}$ Adjusted soft weight of negative pairs
$\mathcal{L}_{NER}$ Classification loss for NER task
$\mathcal{L}_{MS}$ Multi-similarity loss
$\mathcal{L}_{RMS}$ Refined multi-similarity loss
$\theta$ Classifier parameter
$\epsilon$ Margin penalty
$\alpha$, $\beta$, $\gamma$, $\rho$, $\tau$ Hyperparameters
$\lambda_{\mathbb{S}}$, $\lambda_{\mathbb{T}}$ Balance term for losses
Table 9: Notation Table
Appendix C Details of Event Embedding Construction.
C.1 Concatenation based Event Embedding
@@
C.2 Sentence-Encoder based Event Embedding
The prompt we use to instruct ChatGPT to generate explanatory templates for events is: give a one-sentence definition of biomedical event type XXX with arguments XXX, XXX…. For instance, we generate a template for the Phosphorylation event with give a one-sentence definition of biomedical event type Phosphorylation with arguments Trigger, Theme:Molecule, Cause:Molecule, Site:Simple chemical. The full list of the used templates is shown in Table 14, 15, 16, 17, and 18.
Appendix D Experiment Details
@@
Appendix E Full Experiment Results
Full evaluation results are reported in Table 13.
Target Tasks CHEMD BC5CDR DrugProt
Direct Transfer 46.18 78.55 68.17
EG(MS) 46.08 78.87 68.00
EG(concat) 47.03 78.89 67.49
EG(sentEnc) 46.91 79.76 67.93
Table 10: F1(%) scores on three target tasks. Performance of our EG method using vanilla MS loss without external knowledge is reported as EG(MS). All the reported scores are averaged over 3 different random seeds.
@@ E.1 Generalization Ability
Event Annotator Target Tasks CHEMD BC5CDR DrugProt
Gold-std Concat 47.03 78.89 67.49
SentEnc 46.91 79.76 67.93
Auto-sys Concat 46.06 79.49 68.31
SentEnc 46.11 79.39 68.04
Table 11: F1(%) scores of our proposed EG methods based on human/machine annotated events. We highlight better scores between Gold-std and Auto-sys annotators under each setting with underlines. All the reported scores are averaged over 3 different random seeds.
@@ E.3 Pseudo Label Usage
Dataset Precision Recall F1
Few-shot 42.73 51.69 46.77
Oracle 73.56 80.03 76.66
Table 12: Comparison of the target-only results for different training set sizes in CHEMDNER.
E.4 Size of Target Set
Since the training documents for the target domain are downsampled to roughly 10% of the original corpus, we compare our downsampled training results (few-shot) with those trained on the entire target set (Oracle). We finetune the BERT model with adapters on the full target dataset CHEMDNER in Table 12. The large gap indicates that the limited training data indeed severely hinders the model’s learning of the task and justifies the need for transferring knowledge from a high-resource source domain.
@@ Appendix G Scientific Artifacts
We list the license used in this paper: PathwayCuration (CC BY-SA 3.0), InfectiousDisease (CC BY-SA 3.0), CancerGenetics (CC BY-SA 3.0), CHEMDNER (CC BY 4.0), BC5CDR (CC BY 4.0), DrugProt (CC BY 4.0), Huggingface Transformers (Apache License 2.0), SapBERT (apache-2.0), S-PubMedBert-MS-MARCO-SCIFACT (apache-2.0), OpenAI (Terms of use888openai.com/policies/terms-of-use). We follow the intended use of all the mentioned existing artifacts in this paper.
Evaluate Datasets CHEMDNER BC5CDR DrugProt
Metrics Precision Recall F1 Precision Recall F1 Precision Recall F1
Few-shot (BERT) 42.73 51.69 46.77 72.44 85.86 78.51 63.80 67.42 65.49
@@
Direct Transfer 42.51 49.70 45.82 66.79 88.79 76.22 60.94 68.70 64.52
EG(concat) 44.75 51.78 48.00 71.49 87.04 78.45 64.45 66.39 65.31
EG(sentEnc) 42.80 50.43 46.25 70.91 86.97 78.09 64.36 67.81 66.03
ED 46.97 52.30 49.48 74.71 83.96 79.06 67.05 67.13 67.08
EG(concat)+ED 43.44 51.54 47.11 72.34 86.09 78.61 66.59 66.13 66.32
EG(sentEnc)+ED 43.34 51.73 47.16 75.66 86.14 80.55 66.15 68.32 67.21
@@
Direct Transfer 41.57 48.30 44.65 74.92 82.51 78.53 61.96 67.81 64.75
EG(concat) 45.26 50.65 47.77 72.28 87.41 79.12 63.76 65.25 64.43
EG(sentEnc) 47.33 50.41 48.80 74.58 86.37 80.04 64.83 63.60 64.20
ED 41.07 46.93 43.78 74.60 85.85 79.80 64.25 69.75 66.63
EG(concat)+ED 43.37 50.26 46.43 73.83 85.48 79.23 62.31 66.87 64.47
EG(sentEnc)+ED 41.50 52.19 46.18 74.70 85.95 79.93 65.06 69.11 67.02
@@
Direct Transfer 45.22 51.80 48.27 72.37 86.56 78.82 62.94 69.61 66.07
EG(concat) 40.59 53.19 46.01 71.44 87.22 78.53 65.31 66.67 65.98
EG(sentEnc) 41.54 52.36 46.33 72.68 86.09 78.82 66.99 65.71 66.30
ED 47.06 50.58 48.75 75.08 86.96 80.57 66.26 73.52 69.68
EG(concat)+ED 45.16 51.72 48.16 73.37 85.06 78.76 65.60 67.67 66.64
EG(sentEnc)+ED 41.89 50.17 45.61 73.81 86.03 79.45 65.59 68.47 66.99
@@
Table 13: Full evaluation results. Pathway Curation, Infectious Diseases and Cancer Genetics are set to be the source domain respectively. Experiment settings are the same as main results reported in Table 3. All the reported scores are averaged over 3 different random seeds.
Type Template
@@
Conversion A specific trigger <Trigger> causes the transformation of a molecule <Theme> into another molecule <Product>.
@@
Phosphorylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the addition of a phosphate group at a particular site <Site>, facilitated by another molecule <Cause>.
@@
Dephosphorylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a phosphate group removed from a particular site <Site>, facilitated by another molecule <Cause>.
@@
Acetylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of an acetyl group at a particular site <Site>, catalyzed by another molecule <Cause>.
@@
Deacetylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> has an acetyl group removed from a particular site <Site>, facilitated by another molecule <Cause>.
@@
Ubiquitination Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the attachment of one or more ubiquitin molecules at a particular site, facilitated by another molecule <Cause>, often involving a simple chemical group <Site> as the site of attachment.
@@
Deubiquitination Indicated by the given trigger <Trigger>, a specific molecule <Theme> has ubiquitin molecules removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
@@
Hydroxylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of a hydroxyl group at a particular site, catalyzed by another molecule <Cause> involving a simple chemical group <Site> as the site of attachment.
@@
Dehydroxylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a hydroxyl group removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
@@
Methylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of a methyl group at a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of attachment.
@@
Demethylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a methyl group removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.